OpenStudy (djbreezy):

what are the x- and y- intercepts of the graph of 5x-4y=-20? A. x-int: 5, Y-int -4 B. X-int: 5, Y-int: 4 C. X-int: -4, Y-int: 5 D. X-int: 4, y-int: -5

OpenStudy (djbreezy):

@Zelda

OpenStudy (mathmale):

5x-4y=-20 y-intercept: Let x=0. Find y. x-intercept: Let y=0. Find x.

OpenStudy (djbreezy):

Idk how to do this @mathmale im terrible at this

OpenStudy (mathmale):

5x-4y=-20 is an equation in both x and y. You could re-write it as 5(x)-4y=-20. Replace that x with '0', please. Type out your result.

OpenStudy (anonymous):

C.